# Stop Looking for “Important Tokens” in Multimodal Language Models: Duplication Matters More

### v1

@@ Appendix A Additional Experiments
Avg. 100% 96.0% 96.9% 96.7% 96.8% 96.8% 94.9% 97.2% 93.9% 81.5%
Table 6: Analysis on how to select the pivot token. This study evaluates pivot tokens, comprising a fixed set of $4$ visual and $4$ text tokens, using various criteria with 128 retained tokens. A-Score denotes the Attention Score. $\spadesuit$ represents selecting token with the highest value as the pivot token. $\heartsuit$ represents selecting the token with the smallest value as the pivot token. For instance, A-Score means selecting the token with the highest value of Attention Score as the pivot token.
A.1 Supplementary Results on Pivot Token Selection
Appendix B Detailed Experiment Settings
B.1 Datasets
@@
POPE. POPE is tailored to assess object hallucination in models. It presents a series of binary questions about the presence of objects in images, using accuracy, recall, precision, and F1 score as metrics. This approach offers a precise evaluation of hallucination levels under different sampling strategies.
ScienceQA. ScienceQA spans a wide array of domains, including natural, language, and social sciences. Questions are hierarchically categorized into 26 topics, 127 categories, and 379 skills, providing a diverse and comprehensive testbed for evaluating multimodal understanding, multi-step reasoning, and interpretability.
@@
B.2 Models
We evaluate DART using various open-source MLLMs. For image understanding tasks, experiments are conducted on the LLaVA family, including LLaVA-1.5-7B [^2] https://huggingface.co/liuhaotian/llava-v1.5-7b (Liu et al., 2024d) and LLaVA-Next-7B [^3] https://huggingface.co/liuhaotian/llava-v1.6-vicuna-7b (Liu et al., 2024c), with the latter used to validate performance on high-resolution images. Furthermore, we validate our method on more advanced models, including Qwen2-VL-7B [^4] https://huggingface.co/Qwen/Qwen2-VL-7B-Instruct (Wang et al., 2024a) and MiniCPM-V-2.6555https://huggingface.co/openbmb/MiniCPM-V-2_6 (Yao et al., 2024b). For video understanding tasks, we use Video-LLaVA (Lin et al., 2023) as the baseline model. following the settings reported in their paper to ensure a fair comparison.
B.3 Baselines
@@
SparseVLM (Zhang et al., 2024b) ranks token importance using cross-modal attention and introduces adaptive sparsity ratios, complemented by a novel token recycling mechanism.
@@
FasterVLM (Zhang et al., 2024a) evaluates token importance via CLS attention in the encoder and performs pruning before interaction with the language model, streamlining the overall process.
@@ B.4 Implementation Details
Appendix C Computational Complexity.
@@
Appendix D Future Works
@@
Appendix E Sparsification Visualization on Different Pivot Token Selection Strategy

### v2

@@ Appendix A Additional Experiments
A.1 Supplementary Results on Pivot Token Selection
A.2 Influence from the Number of Pivot Tokens
Beyond the investigation of pivot token numbers on MME and TextVQA in §5.3, we conduct additional experiments on several representative visual benchmarks to further support our insight. Figure 8 illustrates that our observations on benchmarks such as POPE and SQA align with those in §5.3—namely, that both insufficient and excessive pivot tokens can lead to suboptimal performance. While an insufficient or excessive number of pivot tokens may result in suboptimal outcomes, our statistical analysis reveals that even the worst-performing settings still match or surpass the performance of existing token pruning approaches. This further demonstrates the superiority of DART.
A.3 More Experimental Results on Larger MLLMs
Method MME POPE GQA TextVQA SQA Avg.
Qwen2-VL-72B Upper Bound, Full Tokens  (100%)
Vanilla 2521 87.4 65.3 82.8 91.6 100%
[…]
FastV (ECCV24) 2376 83.8 62.5 81.5 87.6 96.0%
DART (Ours) 2511 85.7 64.2 82.1 90.9 98.9%
[…]
FastV (ECCV24) 2219 81.1 59.2 79.6 85.1 92.1%
DART (Ours) 2496 83.8 62.5 80.4 88.1 96.8%
[…]
FastV (ECCV24) 2089 78.7 55.7 75.4 83.3 88.0%
DART (Ours) 2350 79.3 59.2 76.6 86.0 92.2%
Table 7: Comparative experiments on Qwen2-VL-72B.
@@ A.3 More Experimental Results on Larger MLLMs
Avg. $100\%$ $96.0\%$ $96.9\%$ $96.7\%$ $96.8\%$ $96.8\%$ $94.9\%$ $97.2\%$ $93.9\%$ $81.5\%$
Table 8: Analysis on how to select the pivot token. This study evaluates pivot tokens, comprising a fixed set of $4$ visual and $4$ text tokens, using various criteria with 128 retained tokens. A-Score denotes the Attention Score. $\spadesuit$ represents selecting token with the highest value as the pivot token. $\heartsuit$ represents selecting the token with the smallest value as the pivot token. For instance, A-Score means selecting the token with the highest value of Attention Score as the pivot token. For the Random pivot token selection strategy, we conducted experiments five times using five different random seeds, and report the corresponding standard deviation to reflect variability.
Figure 8: Impact of the number of pivot tokens on performance across additional visual benchmarks. All experiments are conducted with a token reduction ratio of 77.8%. It is noteworthy that even under relatively extreme numbers of pivot tokens, our worst-case performance still matches or surpasses that of existing token pruning methods.
While prior experiments primarily focused on models with 7B parameters, we further validate the effectiveness and robustness of DART on substantially larger models, including LLaVA-v1.5-13B [^3] https://huggingface.co/liuhaotian/llava-v1.5-13b and Qwen2-VL-72B [^4] https://huggingface.co/Qwen/Qwen2-VL-72B-Instruct. Our results demonstrate that DART consistently outperforms prior token pruning methods such as FastV (Chen et al., 2024) and SparseVLM (Zhang et al., 2024c) across various pruning ratios and downstream tasks, while maintaining near-Vanilla performance.
As shown in Table 9, on LLaVA-1.5-13B with an 88.9% pruning ratio, DART achieves 94.7% average performance, significantly outperforming SparseVLM (79.7%) and FastV (81.0%). Similarly, on Qwen2-VL-72B, DART reaches 92.2% under the same pruning ratio, surpassing FastV (88.0%) (Table 7). At a moderate 66.7% pruning ratio, DART retains 99.5% and 98.9% accuracy on LLaVA-1.5-13B and Qwen2-VL-72B, respectively, with minimal degradation.
DART also excels on specific tasks, achieving 60.9 GQA on LLaVA-1.5-13B at 77.8% pruning and 90.9 ScienceQA on Qwen2-VL-72B at 66.7%, both outperforming FastV. These results demonstrate DART ’s scalability and its ability to balance compression and performance in large MLLMs.
@@
LLaVA-1.5-13B Upper Bound, 576 Tokens  (100%)
Vanilla 63.3 68.9 62.3 1818 85.9 72.8 61.3 56.6 100%
[…]
FastV (ECCV24) 59.1 54.0 51.2 1641 82.3 56.4 51.6 56.9 87.8%
SparseVLM (ICML25) 58.7 67.4 61.0 1768 82.2 73.1 45.4 56.5 94.5%
DART (Ours) 62.1 68.2 61.4 1855 84.0 73.6 60.2 57.3 99.5%
[…]
FastV (ECCV24) 57.7 57.9 48.8 1673 79.3 57.0 56.0 55.3 88.2%
SparseVLM (ICML25) 57.9 65.8 55.8 1774 81.1 69.9 49.9 56.3 93.2%
DART (Ours) 60.9 67.4 60.7 1839 81.8 74.3 59.0 57.3 98.5%
[…]
FastV (ECCV24) 53.7 50.9 42.1 1567 69.3 56.8 47.1 56.7 81.0%
SparseVLM (ICML25) 50.6 61.3 54.8 1402 65.0 69.0 22.7 54.5 79.7%
DART (Ours) 57.1 65.4 59.3 1722 75.4 74.1 55.9 57.4 94.7%
Table 9: Comparative experiments on LLaVA-1.5-13B. In all experiments for DART, tokens are pruned after the second layer with $8$ pivot tokens. The pivot tokens are selected based on the maximum K-norm.
Appendix B Extensions to Other Scenarios
B.1 Exploring the Effectiveness of DART in Audio Modalities
Method FLEURs $\downarrow$ LibriSpeech $\downarrow$ Avg. $\downarrow$
Phi-4-Multimodal-Instruct Upper Bound, Full Audio Tokens  (100%)
    Vanilla 3.49 6.40 4.95
[…]
    + Random 8.15 25.23 16.69
    + FastV (ECCV24) 19.82 27.90 23.86
    + DART (Ours) 5.05 6.95 6.00
[…]
    + Random 13.18 39.42 26.3
    + FastV (ECCV24) 34.10 51.60 42.85
    + DART (Ours) 5.84 11.64 8.74
[…]
    + Random 37.57 76.85 57.21
    + FastV (ECCV24) 180.0 88.38 134.19
    + DART (Ours) 18.93 49.13 34.03
Table 10: Comparative experiments on Automatic Speech Recognition tasks. In all experiments for DART, tokens are pruned after the 2nd layer with $8$ pivot tokens. The pivot tokens are selected based on the maximum K-norm. The evaluation metric is Word Error Rate (WER).
In recent years, the integration of audio as a core modality (Abouelenin et al., 2025; Team, 2024; Chu et al., 2024) within Multimodal Large Language Models (MLLMs) has garnered increasing attention. As these models evolve to handle complex, real-world tasks that span language, vision, and sound, the ability to effectively process spoken language becomes crucial. Audio understanding, particularly in the form of automatic speech recognition (ASR), plays a foundational role in applications such as virtual assistants, transcription services, voice-controlled systems, and multimodal reasoning agents. Therefore, beyond the widely explored domains of image and video understanding in the visual modality, we further extend our investigation to evaluate the effectiveness of our proposed method on tasks within the audio modality. To conduct our study, we select Phi-4-Multimodal-Instruct [^5] https://huggingface.co/microsoft/Phi-4-multimodal-instruct, an MLLM with strong audio modality capabilities, and evaluate it on two representative speech benchmarks: FLEURs-en (Conneau et al., 2023) and LibriSpeech-long (Park et al., 2024). As demonstrated in Table 10, our proposed method DART consistently outperforms baseline approaches under varying token reduction ratios on both FLEURs-en and LibriSpeech-long benchmarks. While random pruning and FastV result in substantial degradation in recognition performance, particularly under higher reduction rates, DART maintains significantly lower Word Error Rates (WER), showcasing its robustness and effectiveness in preserving critical audio information even with limited token usage.
B.2 Enhancing VLA Efficiency with DART
SIMPLER Method Retained Tokens PickCan MoveNear Drawer DrawerApple Average FLOPs $\downarrow$ Speedup $\uparrow$
Visual Matching CogACT 256 91.3% 85.0% 71.8% 50.9% 74.8% 100.0% 1.00$\times$
Random Dropping 112 9.7% 20.4% 53.5% 0.0% 20.9% 58.5% 1.20$\times$
FastV 56 92.6% 81.4% 69.8% 52.4% 74.1% 42.0% 1.21$\times$
VLA-Cache - 92.0% 83.3% 70.5% 51.6% 74.4% 80.1% 1.38$\times$
DART 56 95.6% 85.8% 69.9% 49.5% 75.2% 44.7% 1.25$\times$
Variant Aggregation CogACT 256 89.6% 80.8% 28.3% 46.6% 61.3% 100.0% 1.00$\times$
Random Dropping 112 4.0% 16.1% 15.6% 0.0% 8.9% 58.5% 1.20$\times$
FastV 56 91.4% 78.6% 27.6% 50.6% 62.1% 42.0% 1.19$\times$
VLA-Cache - 91.7% 79.3% 32.5% 45.8% 62.3% 82.6% 1.37$\times$
DART 56 92.4% 77.0% 35.9% 52.4% 64.4% 44.7% 1.25$\times$
Table 11: Performance of DART on the CogACT versus the other baselines in the SIMPLER environment. Random Dropping denotes a method involving the random retention of visual tokens.
Building on recent progress in multimodal understanding from vision-language models (Awadalla et al., 2023; Li et al., 2022; Radford et al., 2021; An et al., 2024; Luo et al., 2024), Vision-Language-Action (VLA) models represent a significant step toward embodied intelligence. Systems such as OpenVLA (Kim et al., 2024), CogACT (Li et al., 2024a), $pi_{0}$(Black et al., 2024), and RT-2(Brohan et al., 2023) seamlessly translate multimodal inputs into executable actions. Leveraging large-scale datasets (Fang et al., 2024; O’Neill et al., 2024), these models have demonstrated impressive capabilities in complex robotic manipulation and reasoning tasks. As a potential pathway toward Artificial General Intelligence (AGI), we place great emphasis on improving the efficiency of VLA models through our approach.
To this end, we employ the SIMPLER environment (Li et al., 2024b), a simulation-based benchmark specifically designed for table-top manipulation to evaluate our method. SIMPLER aims to closely mirror real-world dynamics observed in robots such as the Google Robot and WidowX, exhibiting strong consistency between simulated and real-world performance. In this setup, the Vision-Language-Action (VLA) model receives 224$\times$224 RGB image observations along with natural language task instructions (e.g., “Pick coke can”) and generates a sequence of actions in 7-DoF Cartesian space. SIMPLER supports two evaluation configurations: Visual Matching, which emphasizes visual fidelity to real-world scenes, and Variant Aggregations, which introduces variability through changes in lighting, background, and surface textures. For the Google Robot, both configurations include the same set of four tasks: Pick coke can; Move near; Open/close drawer and Open top drawer and place apple. Performance is assessed using success rate as the evaluation metric.
As shown in Table 11, DART demonstrates superior performance compared to other baseline methods in the SIMPLER environment. With only 56 retained visual tokens, DART achieves the highest average success rates of 75.2% and 64.4% in Visual Matching and Variant Aggregation, respectively, outperforming Random Dropping (Wen et al., 2025), FastV (Chen et al., 2024), VLA-Cache (Xu et al., 2025), and even vanilla CogACT (Li et al., 2024a). Moreover, DART significantly reduces computational cost, achieving the lower FLOPs (44.7%), which corresponds to a speedup of 1.25$\times$ compared to the CogACT. These results highlight DART ’s efficiency in maintaining high task performance while substantially reducing computational demands.
Appendix C Detailed Experiment Settings
C.1 Datasets
[…]
C.1.1 Image Understanding
@@
POPE. POPE is tailored to assess object hallucination. It presents a series of binary questions about the presence of objects in images, using accuracy, recall, precision, and F1 score as metrics. This approach offers a precise evaluation of hallucination levels under different sampling strategies.
ScienceQA. ScienceQA spans a wide array of domains, including natural, language, and social sciences. Questions are hierarchically categorized into 26 topics, 127 categories, and 379 skills, providing a diverse and comprehensive testbed for evaluating multimodal understanding, multi-step reasoning, and interoperability.
@@
VizWiz. VizWiz is a visual benchmark designed to assist visually impaired individuals. It contains real-world images captured by blind users, paired with questions they ask about the images. The dataset includes 20,523 training, 4,319 validation, and 8,000 test image-question pairs, with each question accompanied by 10 human-annotated answers. VizWiz challenges models to answer questions accurately or determine if a question is answerable, focusing on practical visual understanding and accessibility.
@@ C.1.2 Video Understanding
C.1.3 Automatic Speech Recognition.
FLEURS. FLEURS is a benchmark for evaluating universal speech representations across 102 languages, built on top of the FLoRes-101 dataset. It contains 12 hours of speech data per language, with parallel speech and text for tasks like ASR, Speech LangID, and cross-modal retrieval.
LibriSpeech-Long. LibriSpeech-Long is a benchmark dataset for long-form speech generation, derived from the original LibriSpeech dataset. It provides 4-minute long continuous speech and corresponding transcripts, enabling the evaluation of long-form speech continuation. This benchmark supports reference-based evaluation for long-form speech tasks and facilitates research in generating coherent and contextually relevant speech over extended durations.
C.1.4 Vision-Language-Action Models Simulation Platform
SIMPLER. SIMPLER is a simulation platform for evaluating real-world robot manipulation policies. It features realistic simulated environments that match common real robot setups (e.g., Google Robot and WidowX) and tasks (e.g., picking and moving objects). By addressing control and visual disparities between simulation and reality, SIMPLER achieves strong correlation with real-world performance, providing a scalable and reproducible evaluation tool.
C.2 Models
We evaluate DART using various open-source MLLMs. For image understanding tasks, experiments are conducted on the LLaVA family, including LLaVA-1.5-7B [^6] https://huggingface.co/liuhaotian/llava-v1.5-7b (Liu et al., 2024d) and LLaVA-Next-7B [^7] https://huggingface.co/liuhaotian/llava-v1.6-vicuna-7b (Liu et al., 2024c), with the latter used to validate performance on high-resolution images. Furthermore, we validate our method on more advanced models, including Qwen2-VL-7B [^8] https://huggingface.co/Qwen/Qwen2-VL-7B-Instruct (Wang et al., 2024a) and MiniCPM-V-2.6999https://huggingface.co/openbmb/MiniCPM-V-2_6 (Yao et al., 2024b). Moreover, to enhance the effectiveness of our proposed method, we also validate DART on larger MLLMs, such as Qwen2-VL-72B and LLaVA-1.5-13B. For video understanding tasks, we use Video-LLaVA (Lin et al., 2023) as the baseline model. following the settings reported in their paper to ensure a fair comparison.
C.3 Baselines
@@
SparseVLM (Zhang et al., 2024c) ranks token importance using cross-modal attention and introduces adaptive sparsity ratios, complemented by a novel token recycling mechanism.
@@
FasterVLM (Zhang et al., 2024b) evaluates token importance via CLS attention in the encoder and performs pruning before interaction with the language model, streamlining the overall process.
@@
FiCoCo (Han et al., 2024) introduces a unified “filter-correlate-compress” paradigm to streamline training-free token reduction in Multimodal Large Language Models (MLLMs).
FitPrune (Ye et al., 2025) proposes a method that generates an efficient token pruning strategy for multi-modal large language models by removing redundant visual tokens. FitPrune is easy to deploy and is designed to meet a predefined computational budget while maintaining model performance.
@@ C.4 Implementation Details
Appendix D Computational Complexity.
@@
Appendix E Future Works
@@
Appendix F Sparsification Visualization on Different Pivot Token Selection Strategy